# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

### v6

@@ 4.2 Arithmetization
Proposition 4.6 (Decidability structure).
@@
$\mathrm{Provable}(\varphi)\equiv\exists p\;\mathrm{check}(p,\varphi)=\mathrm{true}$ (4.83)
@@ 4.3 Arithmetic Representation of Proof Data
Definition 4.7 (Proof witness assignment).
@@
$a=(s,n,t,j_{0},k_{0},\ldots,j_{n-1},k_{n-1},a_{0},m_{0},\ldots,a_{n-1},m_{n-1},\ldots)$ (4.84)
where:
  1. (i)
    $s\in\mathbb{N}$ is the carryless list code for the proof sequence,
  2. (ii)
    $n\in\mathbb{N}$ is the proof length,
  3. (iii)
    $t\in\mathbb{N}$ is the target formula code,
  4. (iv)
    $j_{i},k_{i}\in\mathbb{N}$ are Modus ponens witness indices for line $i$,
  5. (v)
    $a_{i},m_{i}\in\{0,1\}$ are justification selector booleans (axiom vs. Modus ponens),
  6. (vi)
    additional auxiliary variables for bounds, line-extraction chains, and intermediate computations.
@@ 4.4 Polynomial Reflection of Basic Tests
Lemma 4.8 (Equality constraint, degree $2$).
@@
$(x-y)^{2}=0.$ (4.85)
Over $\mathbb{N}$, $(x-y)^{2}=0$ implies xy=0x-y=0, hence $x=y$. Conversely, $x=y$ implies $(x-y)^{2}=0$ immediately. ∎
Lemma 4.9 (Boolean constraint, degree $2$).
@@
b(1b)=0.b(1-b)=0. (4.86)
If $b=0$ or $b=1$, then b(1b)=0b(1-b)=0. Conversely, b(1b)=0b(1-b)=0 forces either $b=0$ or $b=1$ over $\mathbb{N}$. ∎
Lemma 4.10 (Conjunction constraint, degree $2$).
@@
$b=b_{1}b_{2}.$ (4.87)
Definition 4.11 (Axiom matching constraint family).
For each line index $i<n$, define a finite family $P_{\mathrm{axiom}}(a,i)$ of polynomial equalities whose simultaneous satisfaction indicates that the formula code $\mathrm{line}(s,i)$ matches one of the axiom schemas. The construction proceeds by pattern matching on the structural encoding of formulas: (i) extract subformula codes via line extraction and Carryless “unpairing”; (ii) enforce structural equality constraints using Lemma 4.8; (iii) check that the pattern matches $\mathsf{K}$, $\mathsf{S}$, or $\mathsf{EFQ}$ by verifying the implication nesting structure.
Proposition 4.12 (Axiom constraint degree).
@@
Pattern matching and structural equality tests involve only arithmetic extraction (bounded carryless line extraction, unpairing) and quadratic equality constraints (Lemma 4.8). No cubic terms are introduced. ∎
Definition 4.13 (Modus ponens constraint family).
@@
$\mathrm{line}(s,k_{i})=\mathrm{enc}_{\mathrm{form}}(\mathrm{line}(s,j_{i})\to\mathrm{line}(s,i)).$ (4.88)
It comprises: (i) Bound constraints enforcing $j_{i}<i$ and $k_{i}<i$ via four-squares representation; (ii) Structural equality constraints expressing that the implication code at index $k_{i}$ equals the pairing of codes at $j_{i}$ and $i$. Each member of $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree $\leq 2$.
Proposition 4.14 (Modus ponens constraint degree).
@@
Bound constraints use quadratic four-square encodings. Structural equality uses quadratic equality tests and the degree-$2$ pairing encoding. No cubic terms appear within any member of the family. ∎
@@ 4.5 Justification Constraint
Definition 4.15 (Justification disjunction).
@@
$a_{i}(1-a_{i})$ $=0,$ (4.89)
$m_{i}(1-m_{i})$ $=0,$ (4.90)
$a_{i}+m_{i}-1$ $=0,$ (4.91)
$a_{i}\cdot Q$ $=0\quad\text{for each }Q\in P_{\mathrm{axiom}}(a,i),$ (4.92)
$m_{i}\cdot Q$ $=0\quad\text{for each }Q\in\mathcal{P}_{\mathrm{MP}}(a,i).$ (4.93)
Constraints (4.89)–(4.90) enforce boolean values. Constraint (4.91) enforces exclusive selection. Constraints (4.92)–(4.93) activate the corresponding verification polynomial.
Proposition 4.16 (Justification constraint degree).
@@
Constraints (4.89)–(4.91) are degree $\leq 2$. Each axiom constraint family member has degree $\leq 2$; activation $a_{i}\cdot Q=0$ raises degree to at most $3$. Each MP constraint is drawn from the family $\mathcal{P}_{\mathrm{MP}}(a,i)$ of degree $\leq 2$ polynomials; activation $m_{i}\cdot Q=0$ again yields degree $\leq 3$. No product of a selector with a degree-33 polynomial is emitted. ∎
@@ 4.6 Target Constraint
Definition 4.17 (Target matching polynomial).
@@
4.7 The Complete Semantic Constraints
Definition 4.18 (Proof constraint system).
@@ 4.7 The Complete Semantic Constraints
    The target constraint (Definition 4.17).
@@
Remark.
The system $\mathcal{C}_{n,t}$ is the pre-aggregation constraint list emitted by $\mathtt{emit\_cubic\_system}$. It consists of finitely many cubic equations, each of degree at most $3$. The subsequent aggregation layer (Section 4.8) collapses this finite conjunction to a single cubic equation without increasing degree.
Theorem 4.19 (Constraint equivalence).
@@
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists a\in\mathbb{N}^{k}\;\bigwedge_{P\in\mathcal{C}_{n,t}}P(a)=0.$ (4.95)
By directions.
  1. Direction 1.
    Forward. Assume $\mathrm{check}(p,\varphi)=\mathrm{true}$. Construct an assignment $a$ by (i) Setting $s$ to the carryless list code of $p$ (Theorem 3.8), (ii) instantiating extraction-chain witnesses so that $\mathrm{line}(s,i)$ equals the $i$-th formula code, (iii) for each line $i$: if justified by axiom, set $a_{i}=1,m_{i}=0$; if justified by Modus ponens with witnesses $j,k$, set $a_{i}=0,m_{i}=1,j_{i}=j,k_{i}=k$, and (iv) setting all remaining auxiliary variables to satisfy bound and equality constraints. By construction, all constraints in $\mathcal{C}_{n,t}$ vanish under $a$.
  2. Direction 2.
    Reverse. Assume $a$ satisfies all constraints in $\mathcal{C}_{n,t}$. Extract the proof sequence via carryless line decoding: $\varphi_{i}=\mathrm{dec}_{\mathrm{form}}(\mathrm{line}(s,i))$ for $i<n$. The justification constraints ensure each line is either an axiom instance or follows by Modus ponens. The target constraint ensures $\varphi_{n-1}=\varphi$. Hence $\mathrm{check}(p,\varphi)=\mathrm{true}$.
@@
Pre-aggregation cubic system $\mathcal{C}_{n,t}$ (no selector activation)
@@
Axiom $A_{i}(\vec{x})\cdot M_{i}(\vec{x})=0$ 3 Quadratic $\times$ linear disj.
@@
Table 1: Degree bookkeeping for the pre-aggregation cubic constraint system $\mathcal{C}_{n,t}$ after eliminating selector-based activation. Axiom and Modus ponens are combined into a single cubic justification equation $A_{i}\cdot M_{i}=0$, where $A_{i}$ is a quadratic axiom-matching obligation and $M_{i}$ is the (linear-in-the-encoding) MP structural equality. No boolean selectors and no exactly-one constraints are used. The subsequent mixed-radix aggregation layer collapses this finite conjunction to a single cubic equation without increasing degree.
Linear equalities are written as xy=0x-y=0 throughout; squaring is not needed for degree accounting and would only raise degrees.
Proposition 4.20 (Global degree bound).
@@
The mechanization certifies the bound syntactically. Each constraint is emitted as an $\mathtt{Expr}$ whose degree is computed by structural recursion in $\mathtt{R00\_\_Degree\_Framework}$. Axiom constraints are degree $\leq 2$, MP family members are degree $\leq 2$, and activation multiplies these by linear selectors, keeping every emitted equation at degree $\leq 3$. Linear equalities (xy=0x-y=0) are included in both channels to maintain consistent aggregation. The assembly theorem
$\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ (4.96)
proves that every emitted polynomial has syntactic degree $\leq 3$, independently of any assignment. Table 1 gives the expanded concrete bookkeeping that explains why degree $3$ suffices. ∎
Lemma 4.21 (AST degree correspondence).
@@
By structural induction on $e$: leaves $\mathtt{Const}$ and $\mathtt{Var}$ have degree 0 and $1$ respectively; constructors $\mathtt{Add},\mathtt{Sub}$ take the maximum of subdegrees; $\mathtt{Mult}$ adds subdegrees. This matches the algebraic definition of total degree. ∎
Intuition.
@@ 4.8 Mixed-Radix Aggregation
Lemma 4.22 (Two-channel decomposition).
@@
$P(\vec{x})=0\quad\Longleftrightarrow\quad A(\vec{x})=B(\vec{x}).$ (4.97)
@@
Definition 4.23 (Capacity constraints).
@@
Lemma 4.24 (Mixed-radix aggregation).
Let $(A_{i}=B_{i})_{i=1}^{m}$ be equalities of nonnegative polynomials satisfying $0\leq A_{i},B_{i}<B$ for a fixed base $B\geq 2$. Then
$\sum_{i=1}^{m}B^{i}\,A_{i}=\sum_{i=1}^{m}B^{i}\,B_{i}\quad\Longleftrightarrow\quad\forall i,\;A_{i}=B_{i}.$ (4.100)
Under the digit bounds, every natural number has a unique representation $\sum_{i=1}^{m}B^{i}d_{i}$ with $d_{i}\in\{0,\ldots,B-1\}$. Apply uniqueness with digits $d_{i}:=A_{i}$ and $d_{i}:=B_{i}$ on the two sides. ∎
Lemma 4.25 (Aggregation correctness).
Given a finite list of cubic constraints $(L_{i}(\vec{x})=0)_{i=1}^{m}$, let $C_{1},\ldots,C_{r}$ be the associated capacity equalities. For each equality (original or capacity) write $E_{k}=A_{k}-B_{k}$ with $0\leq A_{k},B_{k}<B$. Define the aggregated polynomial
$\Phi(\vec{x},\vec{y}):=\sum_{k=1}^{m+r}B^{k}\,A_{k}(\vec{x},\vec{y})\;-\;\sum_{k=1}^{m+r}B^{k}\,B_{k}(\vec{x},\vec{y}).$ (4.101)
Then the mixed-radix construction outputs the single cubic equation $\Phi(\vec{x},\vec{y})=0$ such that
$\exists\,\vec{x}\;\bigwedge_{i=1}^{m}L_{i}(\vec{x})=0\quad\Longleftrightarrow\quad\exists\,\vec{x},\vec{y}\;\;\Phi(\vec{x},\vec{y})=0,$ (4.102)
where $\vec{y}$ comprises the slack variables for capacity bounds. Moreover, $\deg\Phi\leq 3$.
Extend the list by adding all capacity equalities; enumerate all equalities as $E_{1},\ldots,E_{m+r}$. Decompose each $E_{k}=A_{k}-B_{k}$ with digit bounds <B<B and form $\Phi$ as in (4.101). If all equalities hold, each $A_{k}=B_{k}$ and $\Phi=0$ follows. Conversely, if $\Phi=0$ then by Lemma 4.24 and the digit bounds we get $A_{k}=B_{k}$ for every $k$, so both the original constraints and the capacity equalities hold. Degree is preserved because only constant weights $B^{k}$ are introduced and capacity equations are cubic. ∎
Example 4.26.
@@
Theorem 4.27 (Proof evaluation via single cubic).
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists\,\vec{z}\in\mathbb{N}^{k^{\prime}}:\Phi_{n,t}(\vec{z})=0,$ (4.106)
where $\Phi_{n,t}$ is the single cubic equation produced by banded aggregation of $\mathcal{C}_{n,t}$, and $\vec{z}$ extends the proof witness $\vec{x}$ by the capacity slack variables.
The forward direction follows from Theorem 4.19 and Lemma 4.25.
For the reverse direction, assume $\Phi_{n,t}(\vec{z})=0$. The single mixed-radix equation aggregates every constraint in $\mathcal{C}_{n,t}$, including all capacity equalities. Mixed-radix uniqueness forces each aggregated equality to hold separately, so in particular the digit bounds $0\leq A_{i},B_{i}<B$ hold, and every channel equality $A_{i}=B_{i}$ holds. Therefore all constraints in $\mathcal{C}_{n,t}$ vanish, and validation follows by Theorem 4.19. ∎
Proposition 4.28 (Degree bound).
@@
The aggregation multiplies channel polynomials only by constant powers $B^{i}$. No squaring of cubic expressions occurs. Capacity equations are degree $\leq 3$ (Definition 4.23), and the principal mixed-radix equality is a linear combination of degree-$\leq 3$ channels. Hence the final single equation has syntactic degree $\leq 3$, as certified in $\mathtt{aggregate\_banded\_degree\_le\_3}$. ∎

### v7

@@ 4.2 Arithmetization
Proposition 4.2.1 (Decidability structure).
@@
$\mathrm{Provable}(\varphi)\equiv\exists p\;\mathrm{check}(p,\varphi)=\mathrm{true}$ (100)
@@ 4.3 Arithmetic Representation of Proof Data
Definition 4.3.1 (Proof witness assignment).
@@
$a=(s,n,t,j_{0},k_{0},\ldots,j_{n-1},k_{n-1},a_{0},m_{0},\ldots,a_{n-1},m_{n-1},\ldots)$ (101)
where (i) $s\in\mathbb{N}$ is the carryless list code for the proof sequence, (ii) $n\in\mathbb{N}$ is the proof length, (iii) $t\in\mathbb{N}$ is the target formula code, (iv) $j_{i},k_{i}\in\mathbb{N}$ are modus ponens witness indices for line $i$, (v) $a_{i},m_{i}\in\{0,1\}$ are justification selector booleans (axiom vs. modus ponens), (vi) additional auxiliary variables for bounds, line-extraction chains, and intermediate computations.
@@ 4.4 Polynomial Reflection of Basic Tests
Lemma 4.4.1 (Equality constraint, degree $2$).
@@
$(x-y)^{2}=0.$ (102)
By correctness. Over $\mathbb{N}$, $(x-y)^{2}=0$ implies xy=0x-y=0, hence $x=y$. Conversely, $x=y$ implies $(x-y)^{2}=0$. ∎
Lemma 4.4.2 (Boolean constraint, degree $2$).
@@
b(1b)=0.b(1-b)=0. (103)
By correctness. If $b=0$ or $b=1$, then b(1b)=0b(1-b)=0. Conversely, b(1b)=0b(1-b)=0 forces either $b=0$ or $b=1$ over $\mathbb{N}$. ∎
Lemma 4.4.3 (Conjunction constraint, degree $2$).
@@
$b=b_{1}b_{2}.$ (104)
Definition 4.4.4 (Axiom matching constraint family).
For each line index $i<n$, define a finite family $P_{\mathrm{axiom}}(a,i)$ of polynomial equalities whose simultaneous satisfaction indicates that the formula code $\mathrm{line}(s,i)$ matches one of the axiom schemas. The construction proceeds by pattern matching on the structural encoding of formulas: (i) extract subformula codes via line extraction and Carryless-unpairing; (ii) enforce structural equality constraints using Lemma 4.4.1; (iii) check that the pattern matches $\mathsf{K}$, $\mathsf{S}$, $\mathsf{EFQ}$, or an $\mathsf{RA}$ non-logical axiom (successor, $\Delta_{0}$-induction, $\mathrm{B}\Sigma_{1}$) by verifying the implication nesting structure and the appropriate schema parameters.
Proposition 4.4.5 (Axiom constraint degree).
@@
By construction. Pattern matching and structural equality tests involve only arithmetic extraction (bounded Carryless-line extraction, unpairing) and quadratic equality constraints (Lemma 4.4.1). No cubic terms are introduced. ∎
Definition 4.4.6 (Modus ponens constraint family).
@@
$\mathrm{line}(s,k_{i})=\mathrm{enc}_{\mathrm{form}}(\mathrm{line}(s,j_{i})\to\mathrm{line}(s,i)).$ (105)
It comprises: (i) Bound constraints enforcing $j_{i}<i$ and $k_{i}<i$ via Four-squares, (ii) Structural equality constraints expressing that the implication code at index $k_{i}$ equals the pairing of codes at $j_{i}$ and $i$. Each member of $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree $\leq 2$.
Proposition 4.4.7 (Modus ponens constraint degree).
@@
Immediate. Bound constraints use quadratic Four-square encodings. Structural equality uses quadratic equality tests and the degree-$2$ pairing encoding. No cubic terms appear within any member of the family. ∎
@@ 4.5 Justification Constraint
Definition 4.5.1 (Justification disjunction).
@@
$a_{i}(1-a_{i})$ $=0,$ (106)
$m_{i}(1-m_{i})$ $=0,$ (107)
$a_{i}+m_{i}-1$ $=0,$ (108)
$a_{i}\cdot Q$ $=0\quad\text{for each }Q\in P_{\mathrm{axiom}}(a,i),$ (109)
$m_{i}\cdot Q$ $=0\quad\text{for each }Q\in\mathcal{P}_{\mathrm{MP}}(a,i).$ (110)
Constraints (106)–(107) enforce boolean values. Constraint (108) enforces exclusive selection. Constraints (109)–(110) activate the corresponding verification polynomial.
Proposition 4.5.2 (Justification constraint degree).
@@
By exhaustion. Constraints (106)–(108) are degree $\leq 2$. Each axiom constraint family member has degree $\leq 2$; activation $a_{i}\cdot Q=0$ raises degree to at most $3$. Each MP constraint is drawn from the family $\mathcal{P}_{\mathrm{MP}}(a,i)$ of degree $\leq 2$ polynomials; activation $m_{i}\cdot Q=0$ again yields degree $\leq 3$. No product of a selector with a degree-$3$ polynomial is emitted. ∎
@@ 4.6 Target Constraint
Definition 4.6.1 (Target matching polynomial).
@@
4.7 Semantic Constraints
Definition 4.7.1 (Proof constraint system).
@@ 4.7 Semantic Constraints
    The target constraint (Definition 4.6.1).
@@
Remark 4.7.2.
The system $\mathcal{C}_{n,t}$ is the pre-aggregation constraint list emitted by $\mathtt{emit\_cubic\_system}$. It consists of finitely many equations, each of degree at most $3$. The subsequent aggregation layer (Section 4.8) collapses this finite conjunction to a single cubic equation without increasing degree.
Theorem 4.7.3 (Constraint equivalence).
@@
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists a\in\mathbb{N}^{k}\;\bigwedge_{P\in\mathcal{C}_{n,t}}P(a)=0.$ (112)
By correctness.
Forward. Assume $\mathrm{check}(p,\varphi)=\mathrm{true}$. Construct an assignment $a$ by (i) Setting $s$ to the Carryless-list code of $p$ (Theorem 3.3.4), (ii) instantiating extraction-chain witnesses so that $\mathrm{line}(s,i)$ equals the $i$-th formula code, (iii) for each line $i$: if justified by axiom, set $a_{i}=1,m_{i}=0$; if justified by modus ponens with witnesses $j,k$, set $a_{i}=0,m_{i}=1,j_{i}=j,k_{i}=k$, and (iv) setting all remaining auxiliary variables to satisfy bound and equality constraints. By construction, all constraints in $\mathcal{C}_{n,t}$ vanish under $a$.
Reverse. Assume $a$ satisfies all constraints in $\mathcal{C}_{n,t}$. Extract the proof sequence via Carryless-line extraction: $\varphi_{i}=\mathrm{dec}_{\mathrm{form}}(\mathrm{line}(s,i))$ for $i<n$. The justification constraints ensure each line is either an axiom instance or follows by modus ponens. The target constraint ensures $\varphi_{n-1}=\varphi$. Hence $\mathrm{check}(p,\varphi)=\mathrm{true}$. ∎
@@
Boolean selectors $a_{i}(1-a_{i})=0,\;m_{i}(1-m_{i})=0$ 2 Forces $a_{i},m_{i}\in\{0,1\}$
Exclusivity $a_{i}+m_{i}-1=0$ 1 Exactly one active
Axiom activation $a_{i}\cdot A_{i}(\vec{x})=0$ 3 Selector $\times$ quadratic
MP activation $m_{i}\cdot M_{i}(\vec{x})=0$ 3 Selector $\times$ quadratic
@@
Table 1: Degree bookkeeping for the pre-aggregation cubic constraint system $\mathcal{C}_{n,t}$. Justification is enforced by selector-activated constraints $a_{i}\cdot A_{i}=0$ and $m_{i}\cdot M_{i}=0$, with exactly-one enforced by $a_{i}+m_{i}=1$. Each activation multiplies a degree-$1$ selector by a degree-$\leq 2$ obligation, yielding degree $\leq 3$. The subsequent mixed-radix aggregation layer collapses this finite conjunction to a single cubic equation without increasing degree.
Proposition 4.7.4 (Global degree bound).
@@
By exhaustion. Each polynomial in $\mathcal{C}_{n,t}$ is obtained from one of the following constructions. Linear extraction and readout constraints such as $t_{i,0}-s=0$ and $f_{i}-h_{i,i}=0$ have degree $1$.
Boolean well-formedness constraints b(1b)=0b(1-b)=0 have degree $2$ (Lemma 4.4.2). Equality tests $(x-y)^{2}=0$ have degree $2$ (Lemma 4.4.1). Zeckendorf digit and non-adjacency constraints are of the form $d-d^{2}=0,$ and $d_{\kappa}d_{\kappa+1}=0,$ hence degree $2$. By Proposition 4.4.5 and 4.4.7, every polynomial in $P_{\mathrm{axiom}}(a,i)$ and $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree at most $2$. Activated obligations are emitted in the form $a_{i}\cdot Q=0,\quad m_{i}\cdot Q=0,$ where $a_{i},m_{i}$ are selector variables and $Q$ is one of the above degree-$\leq 2$ polynomials. Since $a_{i}$ and $m_{i}$ are variables of degree $1$, each activated constraint has total degree at most $1+2=3$. No constraint multiplies a selector with a polynomial of degree $3$. See Table 1. The target condition
$(\mathrm{line}(s,n-1)-t)^{2}=0$ (113)
has degree $2$ (Definition 4.6.1). These cases exhaust all emitted constraints in $\mathcal{C}_{n,t}$. ∎
Lemma 4.7.5 (AST degree correspondence).
@@
By induction on the structure of $e$: leaves $\mathtt{Const}$ and $\mathtt{Var}$ have degree 0 and $1$ respectively; constructors $\mathtt{Add},\mathtt{Sub}$ take the maximum of subdegrees; $\mathtt{Mult}$ adds subdegrees. This matches the algebraic definition of total degree. ∎
Remark 4.7.6.
@@ 4.8 Mixed-Radix Aggregation
Lemma 4.8.2 (Two-channel decomposition).
@@
$P(\vec{x})=0\quad\Longleftrightarrow\quad A(\vec{x})=B(\vec{x}).$ (114)
@@
Definition 4.8.3 (Capacity constraints).
@@
Lemma 4.8.4 (Mixed-radix aggregation).
Let $(A_{i}=B_{i})_{i=1}^{m}$ be equalities of nonnegative polynomials satisfying
$0\leq A_{i},B_{i}<B$ (117)
for a fixed base $B\geq 2$. Then
$\sum_{i=1}^{m}B^{i}\,A_{i}=\sum_{i=1}^{m}B^{i}\,B_{i}\quad\Longleftrightarrow\quad\forall i,\;A_{i}=B_{i}.$ (118)
By correctness. Under the digit bounds, every natural number has a unique representation
$\sum_{i=1}^{m}B^{i}d_{i},\quad d_{i}\in\{0,\ldots,B-1\},$ (119)
so uniqueness with digits $d_{i}:=A_{i}$ and $d_{i}:=B_{i}$ is applied on the two sides. ∎
Lemma 4.8.5 (Self-bounding aggregation correctness).
Given a finite list of cubic constraints
$(L_{i}(\vec{x})=0)_{i=1}^{m}$ (120)
over $\mathbb{N}$, (Lemma 4.8.2) decompose each
$L_{i}=A_{i}-B_{i}\quad\text{with}\quad A_{i},B_{i}\geq 0.$ (121)
Let $B\geq 2$ be a base determined syntactically from the number of channels $m$ (e.g. $B:=F_{K+1}$ where $K$ is computed from the constraint count), and include the capacity equalities $C_{1},\ldots,C_{r}$ of Definition 4.8.3 for every channel value $A_{i},B_{i}$, enforcing
$B-1-A_{i}(\vec{x})=\textstyle\sum_{j=1}^{4}u_{i,j}^{2},\quad B-1-B_{i}(\vec{x})=\textstyle\sum_{j=1}^{4}v_{i,j}^{2},$ (122)
so that any satisfying assignment automatically satisfies the digit bounds
$0\leq A_{i},B_{i}<B.$ (123)
Enumerate all equalities (original and capacity) as $E_{1},\ldots,E_{m+r}$ and define
$\Phi(\vec{x},\vec{s}):=\sum_{k=1}^{m+r}B^{k}\,A_{k}(\vec{x},\vec{s})\;-\;\sum_{k=1}^{m+r}B^{k}\,B_{k}(\vec{x},\vec{s}).$ (124)
Then, quantifying over all of $\mathbb{N}$ (no external representation bound):
$\exists\,\vec{x}\;\bigwedge_{i=1}^{m}L_{i}(\vec{x})=0\quad\Longleftrightarrow\quad\exists\,\vec{x},\vec{s}\;\;\Phi(\vec{x},\vec{s})=0,$ (125)
where $\vec{s}$ comprises the slack variables for capacity bounds. Moreover, $\deg\Phi\leq 3$.
By correctness.
Forward. Suppose $\vec{x}\in\mathbb{N}^{k}$ satisfies all $L_{i}(\vec{x})=0$. Then each original channel satisfies $A_{i}(\vec{x})=B_{i}(\vec{x})$. For the capacity constraints to be satisfiable it suffices that $A_{i}(\vec{x})<B$ and $B_{i}(\vec{x})<B$; but if $L_{i}(\vec{x})=0$ then $A_{i}(\vec{x})=B_{i}(\vec{x})$, and the compiler emits the base $B$ large enough that every channel value at a satisfying assignment lies in $\{0,\ldots,B-1\}$ (this is a syntactic property of the emitted system, not an assumption on an external bound). By Lagrange’s Four-square Theorem, slack variables $\vec{s}_{i}$ exist satisfying (122), and the capacity channel values equal B1<BB-1<B. All $m+r$ channels have values <B<B, so $\Phi=0$ by construction.
Reverse. Suppose $\Phi(\vec{x},\vec{s})=0$ with $\vec{x},\vec{s}\in\mathbb{N}$. The capacity equations (122) are among the aggregated channels; their satisfaction forces $0\leq A_{i}(\vec{x}),B_{i}(\vec{x})<B$ for every original channel. With the digit bounds in hand, Lemma 4.8.4 (mixed-radix uniqueness) forces each original-channel difference to vanish, hence all $L_{i}(\vec{x})=0$.
Degree is preserved: only constant weights $B^{k}$ are introduced and capacity equations are degree $\leq 3$ (Definition 4.8.3). ∎
Example 4.8.6.
@@
Theorem 4.8.8 (Proof evaluation via single cubic).
Let $\Phi_{n,t}$ be the single cubic equation produced by self-bounding aggregation of $\mathcal{C}_{n,t}$ (Lemma 4.8.5) with base $B(n,t)$ computed from the instance structure (number of channels and bands). The emitted system includes capacity constraints (122) that enforce digit bounds for every channel. For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$:
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists\,\vec{z}\in\mathbb{N}^{k^{\prime}}\;:\;\Phi_{n,t}(\vec{z})=0,$ (129)
where $\vec{z}$ comprises the proof witness variables $\vec{x}$ and the capacity slack variables $\vec{s}$. No external representation bound appears; the digit bounds needed for mixed-radix uniqueness are enforced by the capacity constraints built into $\Phi_{n,t}$.
By correctness.
Forward. From Theorem 4.7.3: the canonical assignment satisfies all constraints in $\mathcal{C}_{n,t}$. The compiler outputs a base $B(n,t)$ large enough that every channel value at a satisfying assignment lies in $\{0,\ldots,B-1\}$; the capacity slacks then exist by Lagrange’s Four-square Theorem. By the forward direction of Lemma 4.8.5, $\Phi_{n,t}(\vec{z})=0$.
Reverse. Assume $\Phi_{n,t}(\vec{z})=0$ with $\vec{z}\in\mathbb{N}^{k^{\prime}}$. The capacity constraints are among the aggregated channels; their satisfaction forces the digit bounds $0\leq A_{i},B_{i}<B(n,t)$. By the reverse direction of Lemma 4.8.5, all constraints in $\mathcal{C}_{n,t}$ vanish, and validation follows by Theorem 4.7.3. ∎
Remark 4.8.9 (Self-bounding aggregation and bounded slices).
In the unbounded $\Sigma^{0}_{1}$-hardness pipeline, no external representation bound $R$ or worst-case magnitude $\mu(R)$ is needed: the compiler outputs an instance with built-in capacity channels that enforce the digit bounds for the fixed base baked into that instance. The base $B(n,t)$ is a syntactic function of the constraint count, computed at compile time.
A bounded regime reappears only in the artifact / slices setting (Section 6), where the base is pinned at artifact-generation time for a given resource parameter $k$; the bounded universality theorem (Theorem 6.1.8) provides the equivalence for each slice. Classical universality (Proposition 6.2.4) follows from $\mathrm{Sat}(t)=\exists k\;\mathrm{Sat}_{B_{k}}(t)$, which absorbs the resource parameter into the witness.
Proposition 4.8.10 (Degree bound).
@@
By construction. The aggregation multiplies channel polynomials only by constant powers $B^{i}$. No squaring of cubic expressions occurs. Capacity equations are degree $\leq 3$ (Definition 4.8.3), and the principal mixed-radix equality is a linear combination of degree-$\leq 3$ channels. Hence the final single equation has syntactic degree $\leq 3$, as certified in $\mathtt{aggregate\_banded\_degree\_le\_3}$. ∎